# ZClassifier: Temperature Tuning and Manifold Approximation via KL Divergence on Logit Space

### v2

@@ 4.2 Latent Structure Analysis
Compared to softmax models, KL-regularized ZClassifier—with ResNet backbone—yields a more informative and diverse latent space. This validates our architectural choice and objective formulation: KL divergence regularization, combined with skip-connected feature extractors, enhances class separability and geometric richness in the logit space.
4.3 Calibration Robustness under Logit Perturbation
To evaluate the stability and calibration robustness of each model, we inject Gaussian noise into the predicted logits at inference time and measure classification accuracy. This simulates logit-level uncertainty and tests how well each model maintains discriminative power when confidence is perturbed.
Figure 3: Accuracy under additive Gaussian logit noise. ResNet ZClassifier remains stable up to STD $\sim$1.0, while Softmax and NoKL variants degrade rapidly.
Table 2: Calibration Accuracy by Logit Noise STD
Noise STD ResNet ZC VGG ZC Softmax NoKL ZC
0.00 0.8408 0.8122 0.8148 0.8326
0.22 0.8391 0.8071 0.7975 0.8157
0.44 0.8331 0.7932 0.7459 0.7563
0.67 0.8259 0.7647 0.6685 0.6917
0.89 0.7988 0.7180 0.5973 0.6066
1.11 0.7513 0.6631 0.5324 0.5398
1.33 0.6808 0.6114 0.4648 0.4725
1.56 0.6126 0.5544 0.4046 0.4276
1.78 0.5467 0.5002 0.3751 0.3902
2.00 0.4970 0.4530 0.3379 0.3549
Observations.
Figure 3 and Table 2 reveal stark differences in robustness:
  • ResNet ZClassifier maintains high accuracy up to noise STD 1.0, degrading gradually. Its performance remains above 50% even at STD 2.0, demonstrating strong resilience to uncertainty.
  • VGG ZClassifier performs reasonably up to STD 0.9, but falls off more steeply than ResNet. This reflects its more fragile latent geometry, as seen in previous sections.
  • Softmax Classifier shows rapid decay after STD 0.5. This is expected, as deterministic logits have no internal regularization or uncertainty modeling to compensate for perturbation.
  • NoKL ZClassifier shows similar degradation to Softmax, confirming that without KL regularization, latent variance is unstructured and does not provide robustness.
Interpretation.
These results suggest that KL-regularized Gaussian modeling of logits plays a critical role in calibration robustness. The structured latent distributions learned by ZClassifier act as an inductive prior that resists small perturbations, unlike softmax logits, which are scale-sensitive and unbounded. The advantage is magnified when combined with skip-connected architectures like ResNet.
4.4 Latent Logit Geometry on Validation Set
To further understand the internal geometry learned by each model, we visualize the latent logit vectors of validation samples from CIFAR-10 using Principal Component Analysis (PCA). We project the high-dimensional class logits into 2D space and color each point by ground-truth class.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 4: PCA projection of latent logit vectors for each model on CIFAR-10 validation set. Each point represents a sample and is colored by its class label.
ResNet ZClassifier.
Figure 4(a) shows tight, radially organized clusters aligned with class identity. Classes are well-separated in logit space and exhibit angular structure, suggesting directional encoding consistent with the KL-regularized Gaussian latent objective.
VGG ZClassifier.
Figure 4(b) demonstrates a similar geometric “fan-shaped” structure, but with more class overlap. Despite the Gaussian loss, VGG fails to fully disentangle logit space. This aligns with the hypothesis that skip connections aid in preserving class-conditional directions.
Softmax Classifier.
In Figure 4(c), logits form a dense, entangled cloud with partial class clusters. Without any latent regularization, softmax logits collapse toward magnitude-based separation. There is no explicit variance modeling, so the geometry lacks class-wise anisotropy.
No-KL ZClassifier.
Figure 4(d) reveals scattered and overlapping clusters with poor separation. Despite the stochastic output, the lack of KL alignment results in unstructured logit space. Classes bleed into one another, highlighting the role of KL loss in shaping semantic directions.
Summary.
These visualizations confirm that: (1) ZClassifier learns more structured latent spaces than softmax-based baselines, (2) KL regularization is essential for geometric disentanglement, and (3) residual architectures like ResNet amplify class separation under Gaussian latent assumptions.
4.5 Linear Discriminability via Fisher LDA
To complement PCA-based geometric analysis, we apply Fisher’s Linear Discriminant Analysis (LDA) to evaluate class-wise linear separability in the latent logit space. LDA projects class logits onto a lower-dimensional space that maximizes inter-class variance while minimizing intra-class spread. This is a useful proxy for testing how well the model organizes its representation under supervised guidance.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 5: Fisher LDA projections of latent logit vectors for four model variants. Each dot corresponds to a validation sample, colored by true class. Clear separation indicates strong class encoding.
ResNet ZClassifier.
In Figure 5(a), class clusters are distinctly separated and linearly aligned. Each class forms a dense, low-variance region, with inter-class gaps reflecting high Fisher scores. This indicates that KL-regularized latent logits align not only with class means but also with class directions in a discriminative space.
VGG ZClassifier.
In contrast, Figure 5(b) shows partial class overlap and fuzzier boundaries. While the general alignment is preserved, classes are less compact, particularly for class 3 and 5. This aligns with previous PCA findings that VGG’s latent structure is flatter and less anisotropic.
Softmax Classifier.
Figure 5(c) shows tightly packed class clusters with significant entanglement. Without stochasticity or regularization, the logits collapse into a small region, reducing linear separability. Class 7, 8, and 9 show substantial overlap, indicating degraded discriminability under linear projections.
No-KL ZClassifier.
Figure 5(d) reveals high overlap and diffuse boundaries. Although stochastic, the lack of KL regularization fails to shape the latent space. Class directions are unstable, and class variances are unregulated, confirming the structural collapse observed in other analyses.
Summary.
Fisher LDA projections confirm that ZClassifier—especially with ResNet backbone—produces logit spaces that are both semantically structured and linearly discriminable. KL divergence not only regularizes variance but also implicitly enforces class-specific directions, facilitating downstream calibration and OOD detection tasks.
4.6 Nonlinear Structure via t-SNE Projection
To visualize global and local semantic structure of the latent logit space, we apply t-distributed Stochastic Neighbor Embedding (t-SNE) to the predicted logits of each model. While PCA and LDA reveal linear separability, t-SNE captures nonlinear manifolds and local neighborhood fidelity across classes.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 6: t-SNE embeddings of latent logit vectors on CIFAR-10 validation set. Colors indicate ground-truth class. Separation, density, and compactness of clusters reflect learned latent geometry.
ResNet ZClassifier.
In Figure 6(a), the logit embeddings form highly structured, radially symmetric clusters with large inter-cluster margins. Each class is clearly defined with tight local grouping, indicating strong preservation of class manifolds in high dimensions. This reflects the effect of KL-regularized Gaussian sampling combined with skip-connected feature extractors.
VGG ZClassifier.
Figure 6(b) shows more fragmented clusters. While most classes remain separated, there are multiple split clusters per class and greater local noise. This suggests that VGG lacks sufficient expressiveness to shape logits as compact distributions in t-SNE space, despite having the same Gaussian loss.
Softmax Classifier.
Figure 6(c) reveals spherical but entangled class regions. Although global structure is evident, local boundaries are less distinct. This is expected as softmax logits are not regularized and only reflect hard decision surfaces rather than probabilistic embeddings.
No-KL ZClassifier.
In Figure 6(d), we observe loosely defined clusters with significant overlap. Interestingly, the stochasticity from Gaussian sampling alone produces some variance, but in the absence of KL regularization, class directions and cluster centers are unaligned—leading to unstable latent encoding.
Summary.
t-SNE results corroborate previous findings: ZClassifier with ResNet backbone induces both global separability and local cohesion in latent logit space. In contrast, removing KL or using less expressive architectures leads to more entangled and diffuse embeddings, even under identical task supervision.
4.7 Gaussianity of Logits via GMM Ellipse Analysis
To evaluate whether the predicted class logits align with class-conditional Gaussian structures, we fit a Gaussian Mixture Model (GMM) to the PCA-projected logit space and visualize per-class covariance ellipses. This provides a geometric test of whether the latent structure exhibits class-specific normality, as expected under our KL-regularized objective.
(a) ResNet ZClassifier
(b) VGG ZClassifier
(c) Softmax Classifier
(d) No-KL ZClassifier
Figure 7: 2D GMM ellipses over PCA-projected logit space. Ellipses represent class-wise Gaussian components (means and covariances). Top and right margins show marginal KDEs along each principal axis.
ResNet ZClassifier.
As shown in Figure 7(a), GMM ellipses for each class are compact, distinct, and oriented in semantically aligned directions. This indicates that the model learns near-isotropic Gaussians with low overlap, consistent with its KL-regularized latent objective. Marginal KDEs show unimodal, well-separated peaks along both principal axes.
VGG ZClassifier.
Figure 7(b) shows elliptical structures with higher anisotropy and partial overlap. Some ellipses are skewed or elongated, suggesting instability in class-wise variance modeling. While still Gaussian-like, the lack of skip connections results in less regularized latent clusters.
Softmax Classifier.
Figure 7(c) reveals spherical but overlapping ellipses. Despite forming separable clusters in PCA, the underlying distributions lack variance regularization, leading to uncontrolled scaling and mixture ambiguity. Marginal distributions are broad and multi-modal, confirming the absence of latent distributional control.
No-KL ZClassifier.
Figure 7(d) illustrates the least structured geometry. Ellipses are large, overlapping, and misaligned, with high inter-class confusion. Although the model samples from Gaussians, the absence of KL divergence means these samples do not converge to meaningful class prototypes. KDEs reflect blurred, entangled distributions.
Summary.
GMM ellipse analysis confirms that KL-regularized ZClassifier (especially with ResNet) best satisfies the latent Gaussian prior assumption. This leads to more coherent uncertainty modeling and downstream improvements in OOD detection and calibration stability.
4.8 Out-of-Distribution Detection with KL Divergence
To assess the robustness of ZClassifier to out-of-distribution (OOD) inputs, we measure the KL divergence between each predicted Gaussian logit distribution and a uniform target distribution. A high KL divergence indicates strong deviation from the learned in-distribution patterns. Figure 8 shows the KL score histograms for in-distribution CIFAR-10 and three types of OOD inputs: SVHN, Gaussian noise, and Uniform noise.
(a) ResNet-18 on SVHN
(b) ResNet-18 on Gaussian
(c) ResNet-18 on Uniform
(d) VGG-11 on SVHN
(e) VGG-11 on Gaussian
(f) VGG-11 on Uniform
(g) NoKL Variant on SVHN
(h) NoKL Variant on Gaussian
(i) NoKL Variant on Uniform
Figure 8: Histograms of KL divergence scores between predicted logits and a uniform distribution. Each plot compares in-distribution CIFAR-10 samples and OOD samples under various ZClassifier variants. Clear separation is achieved with KL-regularized Gaussian modeling.
Quantitative Evaluation.
We report standard OOD detection metrics—AUROC, AUPR, and FPR@95—in Table 3. These metrics evaluate the model’s ability to distinguish between in-distribution (CIFAR-10) and OOD samples (SVHN, Gaussian noise, Uniform noise) based on KL divergence scores computed from the predicted Gaussian logits.
The ResNet-based ZClassifier demonstrates exceptional discriminative capability across all three OOD types. It achieves an AUROC of at least 0.999, an AUPR above 0.9994, and a false positive rate at 95% recall (FPR@95) effectively zero. This indicates that the KL divergence values for in- and out-distribution samples are almost completely separable. We attribute this performance to two factors: (1) the expressiveness of the ResNet backbone, which captures a structured logit geometry; and (2) the KL regularization term, which explicitly enforces parametric uncertainty and separation across class-conditional distributions.
In contrast, the NoKL variant of the ZClassifier collapses in all OOD scenarios. Despite having the same backbone and Gaussian logit structure, the absence of the KL term leads to degenerate behavior. The AUROC falls below 0.03, and the FPR@95 reaches 1.0000—a complete failure in distinguishing OOD inputs. This suggests that the model no longer encodes meaningful class uncertainty in its latent space. Instead, logits become overconfident and collapse toward uninformative directions, resulting in overlapping distributions for in- and out-of-distribution data.
The VGG-based ZClassifier exhibits intermediate performance. On Gaussian and Uniform noise, the model attains AUROC scores of 0.95 and 0.93, respectively, which are competitive. However, the model degrades significantly on the SVHN dataset (AUROC: 0.83, AUPR: 0.56, FPR@95: 0.41). This result may reflect two aspects: the VGG backbone’s weaker inductive bias compared to ResNet and a latent space geometry that is less robust to natural image shifts. The SVHN dataset, although semantically different from CIFAR-10, contains well-structured visual content that is more difficult to reject via density-only cues, unlike pure noise.
In summary, the results highlight that (1) KL regularization is essential for OOD separability in Gaussian logit models, (2) ResNet-18 backbones induce more reliable uncertainty-aware decision boundaries, and (3) model expressiveness and regularization jointly determine the fidelity of probabilistic latent structure necessary for robust detection.
Table 3: OOD Detection Performance via KL Divergence (Higher AUROC/AUPR, Lower FPR@95 are better).
Model OOD Type AUROC AUPR FPR@95
ResNet SVHN 0.9994 0.9994 0.0000
ResNet Gaussian Noise 0.9997 0.9998 0.0000
ResNet Uniform Noise 0.9992 0.9996 0.0000
VGG SVHN 0.8333 0.5573 0.4114
VGG Gaussian Noise 0.9528 0.9511 0.2152
VGG Uniform Noise 0.9298 0.9455 0.4424
NoKL SVHN 0.0122 0.1540 1.0000
NoKL Gaussian Noise 0.0208 0.3076 1.0000
NoKL Uniform Noise 0.0184 0.3093 1.0000
Classification Stability.
We also evaluate OOD detection from a binary classification perspective: given a test sample, the task is to determine whether it originates from the in-distribution (CIFAR-10) or from one of the OOD sources (SVHN, Gaussian noise, Uniform noise), using KL divergence as the decision score. We apply a fixed threshold—optimized on a held-out subset—for each model.
The ResNet-based ZClassifier consistently achieves high binary detection accuracy across all OOD types:
  • SVHN: 98.6% total accuracy (InD recall = 0.95, OOD recall = 1.00)
  • Gaussian noise: 97.0% total accuracy
  • Uniform noise: 97.0% total accuracy
These results indicate that the model maintains stable decision boundaries in the presence of both natural and synthetic domain shifts. Notably, the predicted KL divergence values for OOD samples are sharply concentrated in high-score regions, facilitating confident rejection without degrading InD classification.
By contrast, the NoKL variant of the ZClassifier collapses under the same evaluation. Regardless of the OOD type, the model fails to distinguish samples, resulting in binary classification accuracy near or below 50%—essentially chance level. In particular:
  • All FPR@95 metrics for NoKL are 1.0000, showing no ability to reject OOD
  • Precision and recall for OOD class are near zero in all reports
This reinforces the idea that the KL divergence term is not merely auxiliary—it is structurally essential for forming separable and meaningful latent distributions under uncertainty.
Implications.
The strong separation of KL score distributions between InD and OOD samples validates our core hypothesis: KL divergence from class prototypes acts as a statistically grounded, discriminative signal for distributional shift.
By modeling each class logit as a Gaussian and enforcing alignment with isotropic class prototypes via KL regularization, the ZClassifier creates a probabilistic latent space in which:
  1. 1.
    In-distribution samples exhibit low KL values due to close alignment with learned class priors.
  2. 2.
    OOD samples yield high KL scores due to misaligned means and inflated variances, even when feature-level similarity is nontrivial (e.g., SVHN digits).
This behavior emerges directly from the model’s forward pass—no auxiliary classifier, energy function, or external uncertainty estimator is required. The KL divergence is a by-product of the model’s generative assumption and can be computed efficiently at inference.
Ultimately, ZClassifier serves as both a classifier and an uncertainty estimator in a unified framework, offering principled, calibrated decision-making under distributional shift.
5 Conclusion
We proposed ZClassifier, a probabilistic classifier that replaces deterministic logits with diagonal Gaussian latent variables, trained via KL-regularized variational sampling. This model unifies uncertainty calibration, logit normalization, and latent structure modeling under a single objective. Unlike softmax-based discriminative models, ZClassifier directly parameterizes per-class distributions, enabling robust detection of out-of-distribution (OOD) samples via KL divergence.
Empirical results demonstrate that our approach achieves:
  • Near-perfect OOD detection across natural (SVHN) and synthetic (Gaussian, Uniform noise) domains with AUROC $\geq$ 0.999 using only KL scores,
  • Improved calibration robustness under logit-level perturbations without requiring post-hoc temperature scaling,
  • Superior latent geometric structure, as confirmed via PCA, LDA, t-SNE, and GMM ellipses—especially when paired with expressive backbones like ResNet.
Ablation studies further highlight that KL regularization is essential to prevent latent collapse, and that architectural choices (e.g., skip connections in ResNet) interact meaningfully with the Gaussian logit assumption to produce linearly and nonlinearly separable latent spaces.
In conclusion, ZClassifier offers a simple yet powerful alternative to ensemble-based uncertainty estimation or generative modeling. It achieves calibrated, uncertainty-aware predictions and reliable OOD detection in a single forward pass, using only a lightweight modification to standard classifier architectures. We believe this framework serves as a general-purpose template for probabilistic classification in real-world, shift-prone settings.
@@ 6.1 Limitations
Our evaluation of ZClassifier has been limited to CIFAR-10, a small-scale dataset with only 10 flat classes. This leaves open questions regarding the model’s scalability to more complex domains. For instance, in tasks with a large number of classes (e.g., CIFAR-100 or ImageNet), the diagonal Gaussian assumption for each class may become insufficient to preserve separation. Likewise, the model has not yet been tested in structured prediction tasks such as semantic segmentation, where each pixel is associated with spatially correlated class predictions. In these cases, the independence assumption between class logits may no longer hold.
Another limitation is ZClassifier’s untested performance in class-imbalanced or long-tailed distributions. While Gaussian modeling may intuitively capture uncertainty for low-frequency classes, there is no guarantee that KL divergence will regularize minority logits effectively without explicit balancing. Finally, the assumption that each class logit distribution is isotropic Gaussian—centered at either $\mathcal{N}(1,1)$ (true class) or $\mathcal{N}(0,1)$ (false class)—may break down in fine-grained or hierarchically structured settings, where feature overlap and semantic correlation between classes are significant Ridnik et al. (2023); Hsu et al. (2019).
@@ 6.2 Future Work
Future work could adopt tree-structured or hyperbolic latent priors Nickel and Kiela (2017); Liu and Nickel (2020), enabling classes to share covariance structure based on semantic hierarchy. For example, a parent-child Gaussian prior could allow fine-grained subclasses to inherit from superclass prototypes, improving expressiveness and calibration in taxonomic datasets.
@@
ZClassifier could be extended to handle long-tailed distributions via KL loss reweighting Cao et al. (2019) or variance-scaling for minority classes Ren et al. (2020). In few-shot regimes, recent work has shown that probabilistic embeddings improve generalization Snell et al. (2017); Zhang et al. (2020), suggesting that ZClassifier may offer strong sample efficiency through its uncertainty-aware latent space.
@@
Replacing the softmax head in ViT Dosovitskiy et al. (2021) or CLIP Radford et al. (2021) with a Gaussian logit head could yield more calibrated large-scale classifiers. Techniques like zero-shot classifier tuning or vision-language pretraining could be combined with our logit modeling to enhance robustness and interpretability Liu et al. (2023); Zhang et al. (2023).
@@
Classifier-guided diffusion sampling Dhariwal and Nichol (2021); Ho and Salimans (2022) is a promising application. ZClassifier can provide both mean-based guidance and variance-aware modulation for more controllable generation. This fits well into emerging frameworks that mix probabilistic and generative signals for improved fidelity and safety.
In sum, future research will aim to generalize ZClassifier’s probabilistic structure to hierarchical, low-data, or open-world scenarios while preserving the interpretability and sampling efficiency enabled by its Gaussian foundation.

### v3

@@ 4.2 CIFAR-100 Results
Across CIFAR-10 and CIFAR-100, KL-regularized ZClassifiers—particularly with a ResNet backbone—consistently deliver the most reliable calibration, robust latent geometry, and near-perfect OOD separation on extreme distribution shifts. The advantage becomes more pronounced in CIFAR-100, where higher class complexity amplifies the benefits of structured Gaussian regularization. Without KL regularization, latent logits collapse into indistinguishable distributions, erasing variance information and rendering OOD detection ineffective. These results establish KL-regularized Gaussian logits as a principled and scalable foundation for building classifiers that remain both accurate and uncertainty-aware under diverse and challenging distribution shifts.
@@ 5 Discussion
Building on the empirical results in Section LABEL:sec:experiments, we now discuss the scope, limitations, and promising extensions of ZClassifier. Our findings confirm that KL-regularized Gaussian logits, especially when paired with residual backbones, achieve strong separation under distribution shift and superior calibration. However, as with most controlled experimental studies, these results must be interpreted in light of the dataset scale, task diversity, and modeling assumptions.
@@ 5.1 Limitations
While our evaluations demonstrate the potential of ZClassifier, they are limited to CIFAR-10—an in-distribution dataset with only 10 flat, well-separated classes. This raises open questions about scalability to richer, more complex domains. In large-scale benchmarks such as CIFAR-100, ImageNet, or domain-specific medical datasets, the per-class diagonal Gaussian assumption may be insufficient to preserve separation in latent space. Under such complexity, Gaussian parameters might fail to capture multi-modal or anisotropic logit distributions, potentially reducing OOD separability and calibration robustness.
The current framework has also not been evaluated in structured prediction tasks (e.g., semantic segmentation, dense pose estimation), where outputs are spatially correlated and class predictions vary smoothly over the image. Here, the independence assumption between class logits is likely to break down, suggesting a need for structured latent priors or conditional covariance modeling.
Another notable gap is performance under class-imbalanced or long-tailed distributions. While Gaussian modeling can intuitively capture higher uncertainty for minority classes, there is no guarantee that a uniform KL penalty will regularize these cases effectively without targeted reweighting or variance scaling. This is critical for real-world recognition, where skewed class frequencies are common.
Finally, our isotropic prior design—centered at $\mathcal{N}(1,1)$ for true classes and $\mathcal{N}(0,1)$ for false classes—may oversimplify latent geometry in fine-grained or hierarchically structured tasks. When feature overlap between classes is substantial, or when strong semantic correlations exist (e.g., taxonomies, product hierarchies), the rigid separation implied by this prior may not reflect the underlying structure Ridnik et al. (2023); Hsu et al. (2019).
@@ 5.2 Future Work
A natural extension is to incorporate hierarchical or hyperbolic latent priors Nickel and Kiela (2017); Liu and Nickel (2020) that explicitly encode semantic structure. Tree-structured Gaussian priors could tie the covariance of fine-grained subclasses to their superclasses, enabling information sharing and improving calibration in taxonomic datasets. Such an approach could align latent geometry with ontology-based similarity.
@@
ZClassifier could be adapted to long-tailed distributions via KL loss reweighting Cao et al. (2019) or dynamic variance scaling for minority classes Ren et al. (2020). These mechanisms may prevent variance collapse for rare categories, preserving OOD detection capability under severe imbalance. In few-shot regimes, probabilistic embeddings can improve generalization Snell et al. (2017); Zhang et al. (2020), suggesting that ZClassifier’s uncertainty-aware latent space may yield strong sample efficiency.
@@
Replacing the softmax head in large-scale models such as ViT Dosovitskiy et al. (2021) or CLIP Radford et al. (2021) with a Gaussian logit head could improve calibration and robustness. When combined with zero-shot classifier tuning or vision-language pretraining, this could enhance semantic alignment and interpretability at scale Liu et al. (2023); Zhang et al. (2023).
@@
Classifier-guided diffusion sampling Dhariwal and Nichol (2021); Ho and Salimans (2022) is a compelling generative use case. Unlike standard classifiers, ZClassifier offers both mean-based guidance and variance-aware modulation, enabling finer control over sampling trajectories. This dual signal could increase fidelity and safety, especially for ambiguous prompts or low-data conditions.
Broader Generalization Goals.
Ultimately, we view ZClassifier as a step toward unifying probabilistic inference and discriminative training in a single framework. Future research will focus on extending Gaussian logit modeling to hierarchical, low-data, and open-world scenarios while retaining the interpretability and sampling efficiency demonstrated in Section LABEL:sec:results. Addressing the outlined limitations could transform ZClassifier from a small-scale prototype into a versatile, uncertainty-aware classification head applicable across modalities, architectures, and learning paradigms.